# Classical and quantum thermodynamics in a non-equilibrium regime: Application to thermostatic Stirling engine

### v2

@@ I Introduction
However, to investigate whether the thermodynamic laws are satisfied even in quantum cases where the subsystem and bath are quantum mechanically entangled, it is necessary to treat the bath in a non-Markovian, non-perturbative, and non-factorized manner so that the thermal equilibrium state of the total system satisfies the energy conservation law, including the SB interaction. Therefore, equations of motion derived using the Markovian or rotational wave (or secular) approximation, such as the Lindblad equation and quantum master equation, can only be applied to high-temperature regions where the subsystem exhibits semiclassical dynamics.Tanimura (2006, 2020, 2014, 2015) Therefore, the validity of theories based on the Markov assumptionVan Vu and Saito (2022) or those based on factorized initial conditions, such as the fluctuation theoremEsposito, Harbola, and Mukamel (2009); Campisi, Hänggi, and Talkner (2011); Sagawa (2023) and stochastic thermodynamicsSchmiedl and Seifert (2007); Seifert (2012); Esposito, Lindenberg, and den Broeck (2010); Esposito and den Broeck (2011); Strasberg (2022); Rao and Esposito (2018); Cockrell and Ford (2022); Soret, Cavina, and Esposito (2022) in the quantum case, should be carefully examined.
@@
Over the past 30 years, several methodologies have been developed to accurately describe the effects of quantum entanglement in condensed systems while maintaining strict energy conservation to satisfy the first law of thermodynamics. Such methods, in order of development, include the hierarchical equations of motion (HEOM), Tanimura and Kubo (1989); Tanimura (1990); Ishizaki and Tanimura (2005); Tanimura (2006, 2020, 2014, 2015) the quasi-adiabatic path integral (QUAPI),Makri (1995); Makri and Makarov (1995a, b); Thorwart, Reimann, and Hänggi (2000); Jadhao and Makri (2008); Makri (2014); Segal, Millis, and Reichman (2010) and multiconfigurational time-dependent Hartree (MCTDH),Meyer, Manthe, and Cederbaum (1990); Manthe, Meyer, and Cederbaum (1992); Wang and Thoss (2007) in historical order. Among these, the HEOM approach is ideal for thermodynamic investigations, not only because it can perform numerically “exact” dynamic simulations but also because it can evaluate the change in energy of the heat bath and the SB interaction separately, even for processes far from equilibrium.Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2020, 2021); Song and Shi (2017); Gong et al. (2022); Latune, Pleasance, and Petruccione (2023); Boettcher et al. (2024); Koyanagi and Tanimura (2022a, b, 2024a, 2024b, 2024c) Note that although not numerically “exact”, several non-perturbative approaches have also been developed specifically for thermodynamic systems.Strasberg et al. (2018); Strasberg and Esposito (2020); Zhou, Li, and Galperin (2024)
@@
Although the difficulty of numerical simulation has been solved, a fundamental difference exists between open quantum dynamics theory, which is based on a first-principles description of kinetic systems from a microscopic perspective, and thermodynamic theory, which is based on a phenomenological description of thermal systems from a macroscopic perspective. For example, in quantum mechanics, observables are defined as expectation values, whereas in thermodynamics, they are described by macroscopic intensive and extensive variables. Furthermore, quantum mechanics is a formalism for time evolution, whereas thermodynamics deals primarily with static and quasi-static states near thermal equilibrium. In this respect, most quantum thermodynamics studies are merely kinetic simulations of open quantum dynamics, and their relation to thermodynamics has not been studied in depth. Thus, the fundamental difference between microscopic quantum mechanics and macroscopic thermodynamics raises many open questions, such as whether the Carnot limit can be violated in a quantum case or the existence of Maxwell’s demon.
@@ II Reflections on motive power of heat
In our previous paper,Koyanagi and Tanimura (2024b) we presented system-specific thermodynamic laws described as a system–bath model on the basis of open quantum dynamics theory. Here, we develop the same laws by introducing several thermodynamic “statements” without going into the details of the system, as in traditional thermodynamic theory. By doing so, we clarify the distinctive features of thermodynamic theory that allow it to treat systems in non-equilibrium regimes.
@@ II.1 Laws of thermodynamics applied to work
where $\Phi_{\rm A}^{\rm qst}$ is the DL Massieu potential Massieu (1869); Callen and equality holds under quasi-static changes in the natural variables $\beta(t)$ and $X_{\rm A}(t)$ as $\Phi_{\rm A}^{\rm qst}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]=\tilde{W}_{\rm A}^{ext}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]$. The above inequality states that for any process occurring between two states, the work done by the outside is minimized if the process is quasi-static (or reversible) because no energy is dissipated in the heat bath. Note that if we define work as being done by the subsystem on the outside, the inequality in Eq.(8) is reversed, and the relationship is called the maximum work principle (on the outside). Although not mathematically rigorous, the derivation of Eq. (8) is described in Appendix A of Ref. Koyanagi and Tanimura, 2024b. Using the TDL transformation (6), we also obtain the Planck potential Planck (1922); Callen in terms of natural variables expressed as $\Xi_{\rm A}^{\rm qst}[\beta^{\rm qst},\tilde{x}^{\rm qst}]$ (see Appendix A). Similar expressions for Planck functions were derived on the basis of stochastic thermodynamics.Esposito and den Broeck (2011); Rao and Esposito (2018); Strasberg (2022) However, our expression was obtained from the well-defined kinetic system and is valid even under non-factorized SB interactions, where bathentanglement plays an essential role. Thus, the value of $\Delta\Phi_{\rm A}^{\rm qst}$ appearing in Eq. (8) is generally different from the value evaluated from stochastic thermodynamics.
@@
which corresponds to the principle of maximum entropy, where $\Lambda_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},{X}_{\rm A}^{\rm qst}]$ is the entropic potential. Because both natural variables are extensive, this potential is fundamental and we call it the Massieu entropy (M-entropy). Using the TDL transformation given by Eq. (46), we also obtain the Planck entropy (P-entropy) defined by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}]$ from $\tilde{Q}_{\rm A}^{int}(t)$ (see Appendix A). Similar to the entropy obtained from the partition function in statistical physics, this entropy is a function of the extensive variable $\tilde{x}^{\rm qst}$.
Here, as the conjugate variable for $\beta(t)$, we chose the internal energy $U_{\rm A}^{\rm qst}$, which is considered a fundamental variable in thermodynamics. As demonstrated in our previous paper,Koyanagi and Tanimura (2024b) enthalpy $H_{\rm A}^{\rm qst}$ could also be chosen as the conjugate variable for $\beta(t)$. In such a case, there are also two entropies, one involving both extensive variables (Clausius entropy or C-entropy) and one intensive variable with respect to external forces (Boltzmann entropy or B-entropy). The B-entropy and M-entropy values are related to the others by Legendre transformations between $U_{\rm A}^{\rm qst}$ and $H_{\rm A}^{\rm qst}$ and the values do not change, although the former includes one intensive variable, whereas the later includes extensive variables only. However, there is no explicit relationship between C-entropy and M-entropy, because the Legendre transformations between two entropies in the $U_{\rm A}^{\rm qst}$ representation and the $H_{\rm A}^{\rm qst}$ representation are different. (See Appendix D in Ref. Koyanagi and Tanimura, 2024b). In both cases, the Massieu and Planck potentials are the same entropic potentials.Guggenheim (1986)
@@ II.2.2 Non-equilibrium Helmholtz–Gibbs potentials
Helmholtz Energy $\frac{d}{dt}F^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $T,x$
Gibbs Energy $\frac{d}{dt}G^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $T,X_{\rm A}^{\rm neq}$ $G^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
Internal Energy $\frac{d}{dt}U^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $S_{\rm A}^{\rm neq},x$ $U^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}+TS^{\rm neq}_{\rm A}$
Enthalpy $\frac{d}{dt}H^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $S_{\rm A}^{\rm neq},X_{\rm A}^{\rm neq}$ $H^{\rm neq}_{\rm A}=U^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
@@ III.1 Thermodynamic system-bath model
While the HEOM approach to thermodynamics treats the SB interaction as part of the main system in spin-boson systems,Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2021, 2020); Koyanagi and Tanimura (2022a, b, 2024a) this Brownian model treats it as part of a heat bath, including the counter term.Koyanagi and Tanimura (2024b, c) Assuming the Ohmic SDF, in the classical and high-temperature semi-classical cases, the dynamics described by this model exhibit a Markovian future, which can be treated in the framework of stochastic thermodynamics, while in the fully quantum case, it is not only non-Markovian, but also non-factorized owing to bathentanglement. In other words, the difference between the classical and quantum results represents a deviation from Markovian thermodynamics that arises when dealing with fully quantum processes.
@@ Acknowledgments
Shoki Koyanagi: Conceptualization (supporting); Formal analysis (lead); Investigation (equal); Software (lead); Writing - original draft (equal). Yoshitaka Tanimura: Conceptualization (lead); Formal analysis (supporting); Investigation (equal); Writing - review and editing (lead).
@@ Appendix B DL non-equilibrium minimum work principle
By introducing a second non-equilibrium state $B$ on the pathway $A\rightarrow 2$ as depicted in Fig. 7, we now discuss the non-equilibrium transition ${A\rightarrow B}$ (red arrow). We now consider the non-equilibrium-to-equilibrium minimum work from $B$ to 2 (light blue curve), expressed as
@@ Appendix C Convexity of non-equilibrium thermodynamic potentials as functions of extensive variables
To obtain stable thermodynamic properties, the thermodynamic potential must be a convex function of the work variables. Even in the non-equilibrium case, we find that convexity holds under certain conditions. Here, we present this in the SB model. Because other potentials can be discussed similarly, we have limited our discussion to the case of non-equilibrium Massieu potential.
@@ Appendix D Inequality between non-equilibrium and quasi-static thermodynamic potentials
where we have assumed that $\beta(t_{1})=\beta_{2}$ and $X_{\rm A}^{\rm neq}(t_{1})=(X_{\rm A}^{\rm qst})_{2}$, and that $\beta_{2}=\beta(t)$ and $\tilde{x}_{2}=\tilde{x}(t)$ for time $t>t_{1}$, allowing the subsystem to relax to the equilibrium state 2 at time $t_{2}$. Because $X_{\rm A}^{\rm neq}(t_{2})=(X_{\rm A}^{\rm qst})_{2}=X_{\rm A}^{\rm neq}(t_{1})$, the right-hand side of Eq. (D) vanishes. Thus, the net DL extensive work is zero, and so the DL minimum work principle reduces to $0\geq(\Phi_{\rm A}^{\rm neq})_{A\rightarrow 2}$, from which it follows that $\Phi_{\rm A}^{\rm qst}(t_{2})\geq\Phi_{\rm A}^{\rm neq}(t_{1})$.

### v3

@@ I Introduction
However, to investigate whether the thermodynamic laws are satisfied even in quantum cases where the subsystem and bath are quantum mechanically entangled, it is necessary to treat the bath in a non-Markovian, non-perturbative, and non-factorized manner so that the thermal equilibrium state of the total system satisfies the energy conservation law, including the SB interaction. Thus, equations of motion derived using the Markovian or rotational wave (or secular) approximation, such as the Lindblad equation and quantum master equation, can only be applied to high-temperature regions where the subsystem exhibits semiclassical dynamics.Tanimura (2006, 2020, 2014, 2015) Therefore, the validity of theories based on the Markov assumptionVan Vu and Saito (2022) or those based on factorized initial conditions, such as the fluctuation theoremEsposito, Harbola, and Mukamel (2009); Campisi, Hänggi, and Talkner (2011); Sagawa (2023) and stochastic thermodynamicsSchmiedl and Seifert (2007); Seifert (2012); Esposito, Lindenberg, and den Broeck (2010); Esposito and den Broeck (2011); Strasberg (2022); Rao and Esposito (2018); Cockrell and Ford (2022); Soret, Cavina, and Esposito (2022) in the quantum case, should be carefully examined.
@@
Over the past 30 years, several methodologies have been developed to accurately describe the effects of quantum entanglement in condensed systems while maintaining strict energy conservation to satisfy the first law of thermodynamics. Such methods include the hierarchical equations of motion (HEOM), Tanimura and Kubo (1989); Tanimura (1990); Ishizaki and Tanimura (2005); Tanimura (2006, 2020, 2014, 2015) the quasi-adiabatic path integral (QUAPI),Makri (1995); Makri and Makarov (1995a, b); Thorwart, Reimann, and Hänggi (2000); Jadhao and Makri (2008); Makri (2014); Segal, Millis, and Reichman (2010) and multiconfigurational time-dependent Hartree (MCTDH),Meyer, Manthe, and Cederbaum (1990); Manthe, Meyer, and Cederbaum (1992); Wang and Thoss (2007) in historical order. Among these, the HEOM approach is ideal for thermodynamic investigations, not only because it can perform numerically “exact” dynamic simulations but also because it can evaluate the change in energy of the heat bath and the SB interaction separately, even for processes far from equilibrium.Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2020, 2021); Song and Shi (2017); Gong et al. (2022); Latune, Pleasance, and Petruccione (2023); Boettcher et al. (2024); Koyanagi and Tanimura (2022a, b, 2024a, 2024b, 2024c) Note that although not numerically “exact”, several non-perturbative approaches have also been developed specifically for thermodynamic systems.Strasberg et al. (2018); Strasberg and Esposito (2020); Zhou, Li, and Galperin (2024)
@@
While the difficulty of numerical simulation has been solved, a fundamental difference exists between open quantum dynamics theory, which is based on a first-principles description of kinetic systems from a microscopic perspective, and thermodynamic theory, which is based on a phenomenological description of thermal systems from a macroscopic perspective. For example, in quantum mechanics, observables are defined as expectation values, whereas in thermodynamics, they are described by macroscopic intensive and extensive variables. Furthermore, quantum mechanics is a formalism for time evolution, whereas thermodynamics deals primarily with static and quasi-static states near thermal equilibrium. In this respect, most quantum thermodynamics studies are merely kinetic simulations of open quantum dynamics, and their relation to thermodynamics has not been studied in depth. Thus, the fundamental difference between microscopic quantum mechanics and macroscopic thermodynamics raises many open questions, such as whether the Carnot limit can be violated in a quantum case or the existence of Maxwell’s demon.
@@ II Reflections on motive power of heat
In our previous paper,Koyanagi and Tanimura (2024b) we presented system-specific thermodynamic laws described as a system-bath model on the basis of open quantum dynamics theory. Here, we develop the same laws by introducing several thermodynamic “statements” without going into the details of the system, as in traditional thermodynamic theory. By doing so, we clarify the distinctive features of thermodynamic theory that allow it to treat systems in non-equilibrium regimes.
@@ II.1 Laws of thermodynamics applied to work
where $\Phi_{\rm A}^{\rm qst}$ is the DL Massieu potential Massieu (1869); Callen (1991) and equality holds under quasi-static changes in the natural variables $\beta(t)$ and $X_{\rm A}(t)$ as $\Phi_{\rm A}^{\rm qst}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]=\tilde{W}_{\rm A}^{ext}[\beta^{\rm qst},X_{\rm A}^{\rm qst}]$. The above inequality states that for any process occurring between two states, the work done by the outside is minimized if the process is quasi-static (or reversible) because no energy is dissipated in the heat bath. Note that if we define work as being done by the subsystem on the outside, the inequality in Eq.(8) is reversed, and the relationship is called the maximum work principle (on the outside). Although not mathematically rigorous, the derivation of Eq. (8) is described in Appendix A of Ref. Koyanagi and Tanimura, 2024b. Using the TDL transformation (6), we also obtain the Planck potential Planck (1922); Callen (1991) in terms of natural variables expressed as $\Xi_{\rm A}^{\rm qst}[\beta^{\rm qst},\tilde{x}^{\rm qst}]$ (see Appendix A). Similar expressions for Planck functions were derived on the basis of stochastic thermodynamics.Esposito and den Broeck (2011); Rao and Esposito (2018); Strasberg (2022) However, our expression was obtained from the well-defined kinetic system and is valid even under non-factorized SB interactions, where bathentanglement plays an essential role. Thus, the value of $\Delta\Phi_{\rm A}^{\rm qst}$ appearing in Eq. (8) is generally different from the value evaluated from stochastic thermodynamics.
@@
which corresponds to the principle of maximum entropy, where $\Lambda_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},{X}_{\rm A}^{\rm qst}]$ is the entropic potential. Because both natural variables are extensive, this potential is fundamental and we call it the Massieu entropy (M-entropy). Using the TDL transformation given by Eq. (46), we also obtain the Planck entropy (P-entropy) defined by $\Gamma_{\rm A}^{\rm qst}[U_{\rm A}^{\rm qst},\tilde{x}^{\rm qst}]$ from $\tilde{Q}_{\rm A}^{int}(t)$ (see Appendix A). Similar to the entropy obtained from the partition function in statistical physics, this entropy is a function of the intensive variable $\tilde{x}^{\rm qst}$.
Here, as the conjugate variable for $\beta(t)$, we chose the internal energy $U_{\rm A}^{\rm qst}$, which is considered a fundamental variable in thermodynamics. As demonstrated in our previous paper,Koyanagi and Tanimura (2024b) enthalpy $H_{\rm A}^{\rm qst}$ could also be chosen as the conjugate variable for $\beta(t)$. In such a case, there are also two entropies, one involving both extensive variables (Clausius entropy or C-entropy) and one intensive variable with respect to external forces (Boltzmann entropy or B-entropy). The B-entropy and M-entropy values are related to the others by Legendre transformations between $U_{\rm A}^{\rm qst}$ and $H_{\rm A}^{\rm qst}$ and the values do not change, although the former includes one intensive variable, whereas the latter includes extensive variables only. However, there is no explicit relationship between C-entropy and M-entropy, because the Legendre transformations between two entropies in the $U_{\rm A}^{\rm qst}$ representation and the $H_{\rm A}^{\rm qst}$ representation are different. (See Appendix D in Ref. Koyanagi and Tanimura, 2024b). The Massieu and Planck potentials are equivalent to the entropy potentials introduced earlier.Guggenheim (1986)
@@ II.2.2 Non-equilibrium Helmholtz–Gibbs potentials
Helmholtz Energy $\frac{d}{dt}F^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $T,X_{\rm A}^{\rm neq}$
Gibbs Energy $\frac{d}{dt}G^{\rm neq}_{\rm A}=-S^{\rm neq}_{\rm A}\frac{d}{dt}T-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $T,x$ $G^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
Internal Energy $\frac{d}{dt}U^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}+x\frac{d}{dt}X_{\rm A}^{\rm neq}+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $S_{\rm A}^{\rm neq},X_{\rm A}^{\rm neq}$ $U^{\rm neq}_{\rm A}=F^{\rm neq}_{\rm A}+TS^{\rm neq}_{\rm A}$
Enthalpy $\frac{d}{dt}H^{\rm neq}_{\rm A}=T\frac{d}{dt}S^{\rm neq}_{\rm A}-X_{\rm A}^{\rm neq}\frac{d}{dt}x+\frac{d}{dt}Q_{\rm A}^{\rm wst}$ $S_{\rm A}^{\rm neq},x$ $H^{\rm neq}_{\rm A}=U^{\rm neq}_{\rm A}-xX_{\rm A}^{\rm neq}$
@@ III.1 Thermodynamic system-bath model
While the HEOM approach to thermodynamics treats the SB interaction as part of the main system in spin-boson systems,Kato and Tanimura (2015, 2016); Sakamoto and Tanimura (2021, 2020); Koyanagi and Tanimura (2022a, b, 2024a) this Brownian model treats it as part of a heat bath, including the counter term.Koyanagi and Tanimura (2024b, c) Assuming the Ohmic SDF, in the classical and high-temperature semi-classical cases, the dynamics described by this model exhibit a Markovian feature, which can be treated in the framework of stochastic thermodynamics, while in the fully quantum case, it is not only non-Markovian, but also non-factorized owing to bathentanglement. In other words, the difference between the classical and quantum results represents a deviation from Markovian thermodynamics that arises when dealing with fully quantum processes.
@@ Acknowledgments
Shoki Koyanagi: Formal analysis (lead); Investigation (equal); Methodology (equal); Software (lead); Writing - original draft (equal). Yoshitaka Tanimura: Conceptualization (lead); Formal analysis (supporting); Funding acquisitiojn (lead); Investigation (equal); Methodology (equal); Writing - review and editing (lead).
@@ Appendix B DL non-equilibrium minimum work principle
By introducing a second non-equilibrium state $B$ on the pathway $A\rightarrow 2$ as depicted in Fig. 7, we now discuss the non-equilibrium transition ${A\rightarrow B}$ (red arrow). We consider the non-equilibrium-to-equilibrium minimum work from $B$ to 2 (light blue curve), expressed as
@@ Appendix C Convexity of non-equilibrium thermodynamic potentials as functions of extensive variables
To obtain stable thermodynamic properties, the potential must be a convex function of the work variables. Even in the non-equilibrium case, we find that convexity holds under certain conditions. Here, we present this in the SB model. Because other potentials can be discussed similarly, we have limited our discussion to the case of non-equilibrium Massieu potential.
@@ Appendix D Inequality between non-equilibrium and quasi-static thermodynamic potentials
where we have assumed that $\beta(t)=\beta_{2}$ for time $t\geq t_{1}$ and $\tilde{x}(t)=\tilde{x}_{2}$ for time $t>t_{1}$ with the the intensive variables $\beta_{2}$ and $\tilde{x}_{2}$, allowing the subsystem to relax to the equilibrium state 2 at time $t_{2}$. The value of $\tilde{x}_{2}$ is set to satisfy the condition $X_{\rm A}^{\rm neq}(t_{2})=X_{\rm A}^{\rm neq}(t_{1})$. Because $X_{\rm A}^{\rm neq}(t_{2})=(X_{\rm A}^{\rm qst})_{2}=X_{\rm A}^{\rm neq}(t_{1})$, the right-hand side of Eq. (D) vanishes. Thus, the net DL extensive work is zero, and so the DL minimum work principle reduces to $0\geq(\Phi_{\rm A}^{\rm neq})_{A\rightarrow 2}$, from which it follows that $\Phi_{\rm A}^{\rm qst}(t_{2})\geq\Phi_{\rm A}^{\rm neq}(t_{1})$.